use the direct variation equation y=1/8x to solve for y when x=3/4
Answer by stanbon(75887)

use the direct variation equation y=1/8x to solve for y when x=3/4
---
y = (1/8)x
---
If x = 3/4,
y = (1/8)(3/4) = 3/32
=========================
Cheers,
Stan H.
===========